OpenStudy (anonymous):

If -3(y-1)=9, find the value of 1/2y

OpenStudy (anonymous):

You need to solve for y first. So for the first step, we distribute -3 to get: -3y+3=9. Next, we subtract 3 from the left hand side and bring it over to the right hand side: -3y+3=9 -3 -3 ______________________ -3y=6 Next, we divide both sides by -3 to get y=-2. Now plugging back in to find (1/2)y, we get (1/2)(-2)=-1.